OpenStudy (anonymous):

simplify the expression: 5x/2 * 1/x

OpenStudy (anonymous):

Step by step answers would be helpful so I can solve the other questions that are similar :/

OpenStudy (dancing_stars23):

\[5x/2 *1/x \] \[a/b *c/d = a*c/b*d\] so 5x*1 =5x 2*x= 2x \[5x/2x\] cancel out the 'x' answer would be 5/2 ; 2.5

OpenStudy (anonymous):

Your x's cancel so you end up with 5/2*1/1= 5/2